A student purchases a computer by obtaining a simple interest loan the computer cost 1500 and the interest rate of the loan is 12% if the loan is to be paid back in weekly installments over two years what is the amount of the interest paid over two years what is the total amount to be paid back and what is the weekly payment amount?
Answer by Theo(13342)

simple interest formula is i = p * r * t.

i is the interest.
p is the principal
r is the interest rate per time period
t is the number of time periods.

in your problem:

p is equal to 1500.
r is equal to 12% per year / 100 = .12 per year / 52 = .0023076923 per week.
t is equal to 2 years * 52 weeks per year = 104 weeks.

if the loan is paid off in 2 year, the total interest paid is 1500 * .0023076923 * 104 = 360.

the weekly payment amount is (1500 + 360) / 104 = 17.88461538.

that's a weekly principal repayment of 1500 / 104 = 14.42307692 plus a weekly interest payment of 360 / 104 = 3.461538462.

add the principal repayment and the interest payment together and you get a total weekly payment of 17.88461538.

round to the nearest penny and the weekly payment is 17.88.

this is simple interest, not to be confused with compound interest, which is a total different animal.